A cruise company would like to estimate the average beer consumption to plan its beer inventory levels on future​ seven-day cruises.​ (The ship certainly​ doesn't want to run out of beer in the middle of the​ ocean!) The average beer consumption over 15 randomly selected​ seven-day cruises was 81,784 bottles with a sample standard deviation of 4,549 bottles.

A. Construct a 99% confidence interval to estimate the average beer consumption per cruise

The 99% confidence interval to estimate the average beer consumption per cruise is from a lower limit of___ bottles to an upper limit of ___ bottles.

B. What's the margin of error?
Answer #1

Solution :

t\alpha /2,df = 2.977

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.977 * (4549 / \sqrt 15)

Margin of error = E = 3497

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

81784 - 3497 < \mu < 81784 + 3497

78287 < \mu < 85281

(78287 , 85281)

Lower limit = 78287

Upper limit = 85281